Grade 12Physical Chemistry

The rate of a chemical reaction doubles for every o 10 C rise of temperature. If the temperature is raised by o 50 C, the rate of the reaction increases by about : ?

sudhanshu
12 Years agoGrade 12

1 Answer

Sunil Kumar FP
11 Years ago
The rate of a chemical reaction doubles for every o 10 C rise of temperature. If the temperature is raised by o 50 C, the rate of the reaction increases by about 32
we have the temperatur coefficient as 2
u^(deltaT/10 =k2/k1
2^5=k2/k1
k2/k1 =32